# Arens extensions of disjointness preserving multilinear operators on Riesz spaces

### v1

@@ 3 Arens extensions of disjointness preserving multilinear operators
(a) Let $y^{***}\in F^{***}$ be a weak*-continuous disjointness preserving functional. Since $y^{***}$ is weak*-continuous, there exists $y^{\ast}\in F^{\ast}$ such that $y^{\ast\ast\ast}=J_{F^{\ast}}(y^{\ast})$. As $y^{\ast\ast\ast}$ is disjointness preserving and and $J_{F}$ is a Riesz homomorphisms, $y^{\ast}=y^{\ast\ast\ast}\circ J_{F}$ is disjointness preserving. Now, Proposition 3.6 yields that $y^{***}\circ AB_{m}^{\rho}(A)=J_{F^{\ast}}(y^{\ast})\circ AB_{m}^{\rho}(A)$ is a disjointness preserving operator.
(b) For each $j\neq i$, take $x_{j}^{\ast\ast}\in E_{j}^{\ast\ast}$. Given $y^{\ast}\in\overline{{\rm span}}\{\varphi\in F^{\ast}:\varphi\text{ is a disjointness preserving}\}$, there exists a sequence $(y_{n}^{\ast})_{n=1}^{\infty}$ in ${\rm span}\{\varphi\in F^{\ast}:\varphi\text{ is a disjointness preserving}\}$ such that $y_{n}^{\ast}\longrightarrow y^{\ast}$, hence $|y_{n}^{\ast}|\longrightarrow|y^{\ast}|$. Moreover, for each $n\in\mathbb{N}$ there exist disjointness preserving linear functionals $\varphi^{1}_{n},\ldots,\varphi_{n}^{k_{n}}\in F^{\ast}$ and scalars $\alpha_{n}^{1},\ldots,\alpha_{n}^{k_{n}}$ such that $y_{n}^{\ast}=\sum\limits_{j=1}^{k_{n}}\alpha_{n}^{j}\varphi_{n}^{j}$. For each $w\in E_{i}^{\ast\ast}$, write
@@
Let $F$ be a Banach lattice such that $F^{\ast}$ has a Schauder basis consisting of disjointness preserving linear functionals. Then all Arens extensions of any $F$-valued disjointness preserving multilinear operator are disjointness preserving.

### v2

@@ 3 Arens extensions of disjointness preserving multilinear operators
(a) Let $y^{***}\in F^{***}$ be a weak*-continuous disjointness preserving functional, and let $y^{\ast}\in F^{\ast}$ be such that $y^{\ast\ast\ast}=J_{F^{\ast}}(y^{\ast})$. As $y^{\ast\ast\ast}$ is disjointness preserving and $J_{F}$ is a Riesz homomorphisms, $y^{\ast}=y^{\ast\ast\ast}\circ J_{F}$ is disjointness preserving. Now, Proposition 3.3 yields that $y^{***}\circ AB_{m}^{\rho}(A)=J_{F^{\ast}}(y^{\ast})\circ AB_{m}^{\rho}(A)$ is a disjointness preserving operator.
(b) For each $j\neq i$, take $x_{j}^{\ast\ast}\in E_{j}^{\ast\ast}$. Given $y^{\ast}\in\overline{{\rm span}}\{\varphi\in F^{\ast}:\varphi\text{ is a disjointness preserving}\}$, there exists a sequence $(y_{n}^{\ast})_{n=1}^{\infty}$ in ${\rm span}\{\varphi\in F^{\ast}:\varphi\text{ is a disjointness preserving}\}$ such that $y_{n}^{\ast}\longrightarrow y^{\ast}$ in $F^{*}$, hence $|y_{n}^{\ast}|\longrightarrow|y^{\ast}|$. Moreover, for each $n\in\mathbb{N}$ there exist disjointness preserving linear functionals $\varphi^{1}_{n},\ldots,\varphi_{n}^{k_{n}}\in F^{\ast}$ and scalars $\alpha_{n}^{1},\ldots,\alpha_{n}^{k_{n}}$ such that $y_{n}^{\ast}=\sum\limits_{j=1}^{k_{n}}\alpha_{n}^{j}\varphi_{n}^{j}$. For each $w\in E_{i}^{\ast\ast}$, write
@@
Let $F$ be a Dedekind complete Banach lattice such that $F^{\ast}$ has a Schauder basis consisting of disjointness preserving linear functionals. Then all Arens extensions of any $F$-valued disjointness preserving regular multilinear operator are disjointness preserving.
@@
From here onwards, we shall use that every Banach lattice is Archimedean [22, Proposition 1.1.8(i)]. Recall that $\theta$ is the permutation of $\{1,2\}$ given by $\theta(1)=2,\theta(2)=1$.
Proposition 3.7.
Let $E_{1},E_{2},F$ be Banach lattices such that $E_{1}^{*}$ is separable and the lattice operations in $F^{**}$ are weak*-sequentially continuous. Let $A\colon E_{1}\times E_{2}\longrightarrow F$ be a regular bilinear operator. If $A$ is disjointness preserving, then the operator
$AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},\cdot)\colon E_{2}^{\ast\ast}\longrightarrow F^{\ast\ast},\,AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},\cdot)(x_{2}^{\ast\ast})=AR_{2}^{\rho}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast}),$
is disjointness preserving for each $x_{1}^{\ast\ast}\in E_{1}^{\ast\ast}$.
Proof.
Let $x_{2}^{\ast\ast},z_{2}^{\ast\ast}\in E_{2}^{\ast\ast}$ be such that $x_{2}^{\ast\ast}\perp z_{2}^{\ast\ast}$. By Proposition 3.1, we have
$AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1}),x_{2}^{\ast\ast})\perp AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1}),z_{2}^{\ast\ast})\text{ for all }x_{1}\in E_{1}.$
Let $x_{1}^{**}\in E_{1}^{**}$ be such that $\|x_{1}^{**}\|=1$. Since $E_{1}^{*}$ is separable, the weak topology of $E_{1}^{**}$ is metrizable on the bounded sets of $E_{1}^{**}$. So, as $J(B_{E_{1}})$ is weak-dense in $B_{E_{1}^{**}}$, there exists a sequence $(x_{1,n})_{n}$ in $B_{E_{1}}$ such that $J(x_{1,n})\xrightarrow{\,\omega^{\ast}}x_{1}^{**}$ in $E_{1}^{**}$. Since the Arens extension $AR_{2}^{\theta}$ is weak continuous in the first variable [13, p. 413], $AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},z_{2}^{\ast\ast})=\omega^{\ast}-\displaystyle\lim_{n}AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),z_{2}^{\ast\ast})$ and $AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast})=\omega^{\ast}-\displaystyle\lim_{n}AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),x_{2}^{\ast\ast})$. Thus,
$|AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast})|$ $\wedge|AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},z_{2}^{\ast\ast})|$
$=\omega^{\ast}-\lim_{n}|AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),x_{2}^{\ast\ast})|\wedge\omega^{\ast}-\lim_{n}|AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),z_{2}^{\ast\ast})|$
$=\omega^{\ast}-\lim_{n}(|AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),x_{2}^{\ast\ast})|\wedge|AR_{2}^{\theta}(A)(J_{E_{1}}(x_{1,n}),z_{2}^{\ast\ast})|)=0.$
Thus far we have $AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast})\perp AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},z_{2}^{\ast\ast})$ for all $x_{1}^{**}\in E_{1}^{**}$ with $\|x_{1}^{**}\|=1$. For $0\neq z_{1}^{**}\in E_{1}^{**}$, $w_{1}^{\ast\ast}:=\frac{z_{1}^{**}}{\|z_{1}^{**}\|}$ satisfies $\|w_{1}^{**}\|=1$ and $AR_{2}^{\theta}(A)(w_{1}^{\ast\ast},x_{2}^{\ast\ast})\perp AR_{2}^{\theta}(A)(w_{1}^{\ast\ast},z_{2}^{\ast\ast})$. It follows that $AR_{2}^{\theta}(A)(z_{1}^{\ast\ast},x_{2}^{\ast\ast})\perp AR_{2}^{\theta}(A)(z_{1}^{\ast\ast},z_{2}^{\ast\ast})$, which proves that the operator $AR_{2}^{\theta}(A)(x_{1}^{\ast\ast},\cdot)$ is disjointness preserving. ∎
Using that the Arens extension $AR_{2}^{id}(A)\colon E_{1}^{\ast\ast}\times E_{2}^{\ast\ast}\longrightarrow F^{\ast\ast}$ is weak continuous in the second variable and Proposition 3.1, the same proof of Proposition 3.7 gives the following.
Proposition 3.8.
Let $E_{1},E_{2},F$ be Banach lattices such that $E_{2}^{*}$ is separable and the lattice operations in $F^{**}$ are weak*-sequentially continuous. Let $A\colon E_{1}\times E_{2}\longrightarrow F$ be a regular bilinear operator. If $A$ is disjointness preserving, then the operator
$AR_{2}^{id}(A)(\cdot,x_{2}^{\ast\ast})\colon E_{1}^{\ast\ast}\longrightarrow F^{\ast\ast},\,AR_{2}^{id}(A)(\cdot,x_{2}^{\ast\ast})(x_{1}^{\ast\ast})=AR_{2}^{id}(A)(x_{1}^{\ast\ast},x_{2}^{\ast\ast}),$
is disjointness preserving for each $x_{2}^{\ast\ast}\in E_{2}^{\ast\ast}$.
Just combine Propositions 3.7 and 3.8 to get the following.
Corollary 3.9.
Let $E_{1},E_{2}$ be Banach lattices with separable duals, and let $F$ be a Banach lattice such that the lattice operations in $F^{**}$ are weak*-sequentially continuous. Let $A\colon E_{1}\times E_{2}\longrightarrow F$ be a regular disjointness preserving bilinear operator whose Arens extensions coincide. Then the (unique) Arens extensions of $A$ is disjointness preserving.
Recall that a Banach space $E$ is Arens regular if every bounded linear operator from $E$ to $E^{\ast}$ is weakly compact. By [26, Theorem 2.2], the Banach space $E$ is Arens regular if and only if the Arens extensions of every continuous bilinear operator from $E^{2}$ to any Banach space are coincident.
Corollary 3.10.
Let $E$ be an Arens regular Banach lattice with separable dual, and let $F$ be a Banach lattice such that the lattice operations in $F^{**}$ are weak*-sequentially continuous. Then, the unique Arens extension of any disjointness preserving regular $m$-linear operator from $E^{2}$ to $F$ is disjointness preserving.
Examples of nonreflexive Arens regular Banach lattices are $c_{0},\ell_{\infty}$ and $C(K)$ for any infinite compact Hausdorff space $K$. In particular, AM-spaces with order unit are Arens regular.
Remark 3.11.
The last four results can be proved, mutatis mutandis, for disjointness preserving regular $m$-linear operators, $m\geq 2$. The reason is that every Arens extension of a continuous $m$-linear operator is weak continuous in one variable. For details, see the proof of [6, Lemma 2.4].